Heat flux and hydrography at the

Main Endeavour vent field

MGG seminar and Final Examination

May 8, 2003

Scott Veirs

scottv@ocean.washington.edu

University of Washington School of Oceanography

Motivation

Plumes in cross flow: Atmospheric ⟶ In the ocean

Plume flux & form implications:

Crustal formation Ocean chemistry Habitat & dispersal Global heat budget

NASA Earth Observatory

– p.2/19

Outline (∼45 minutes)

(10) The Flow Mow study

(10) Current results

(5) Model results

(10) Heat flux results

(5) Conclusions

(5) Acknowledgements

Flow Mow study site: Endeavour segment

Orientation: Start: 47^{∘}54.5’N ∼300 m relief; crest ∼2100 m Valley 10 km long, 1 km wide, 100 m deep Saddle ∼2170 m

Vent fields ∼2 km spacing?

– p.4/19

Orientation: Start: 47^{∘}54.5’N ∼300 m relief; crest ∼2100 m Valley 10 km long, 1 km wide, 100 m deep Saddle ∼2170 m

Vent fields ∼2 km spacing?

– p.4/19

Orientation: Start: 47^{∘}54.5’N ∼300 m relief; crest ∼2100 m Valley 10 km long, 1 km wide, 100 m deep Saddle ∼2170 m

Vent fields ∼2 km spacing?

– p.4/19

Orientation: Start: 47^{∘}54.5’N ∼300 m relief; crest ∼2100 m Valley 10 km long, 1 km wide, 100 m deep Saddle ∼2170 m

Vent fields ∼2 km spacing?

– p.4/19

Orientation: Start: 47^{∘}54.5’N ∼300 m relief; crest ∼2100 m Valley 10 km long, 1 km wide, 100 m deep Saddle ∼2170 m Vent fields ∼2 km spacing?

– p.4/19

Orientation: Start: 47^{∘}54.5’N ∼300 m relief; crest ∼2100 m Valley 10 km long, 1 km wide, 100 m deep Saddle ∼2170 m Vent fields ∼2 km spacing?

– p.4/19

Main Endeavour vent field (MEF)

Geology mapped very well

110 focused “smokers”

Diffuse flow

300×700 m control volume

Max rise height ∼300 m

MEF heat flux puzzles

Power loss per km of ridge: H_{f} + H_{d} ≃ H_{R} = Q_{R}(L_{r} + ∫_{1200}^{200} c_{r}(T)dT) = 42 MW/km ⟹ 84 MW MEF long-term mean

Study  Estimate [MW] Heat flux
Ginster et al., 1994  615±120  H_{f}
Thomson et al., 1992  2500±1525  H_{p}
Baker & Massoth, 1987  4250±2750  H_{p}
Schultz et al, 1992  9000±760  H_{d}

– p.6/19

The “Flow Mow” control volume approach

Net heat flux via advection through A:

H = ∫_{A} ρc_{p}θu · n̂dA ≃ ρc_{p} ∑_{i=1}^{N} Δ_{S}θ_{i}u_{i}ΔA_{i} ≃ ρc_{p}\overline{Δ_{S}θ}\overline{u}A

Instruments: CTD & ABE

– p.8/19

Instruments: Current meters

– p.9/19

Mean flow

– p.10/19

Oscillatory flow over ridge

Δx_{c} ∼2.2 km above ridge ∼1.3 km in valley

Δx\overline{u} in 12 hr: ∼2.2 km at 5 cm/s ∼0.5 km at 1 cm/s

Animation

Figure courtesy Rick Thomson, IOS Canada

– p.11/19

Oscillatory flow in valley

Animation

– p.12/19

Hydrography and heat flux in the valley

H_{h} = \overline{H}_{N} + \overline{H}_{S} = ρc_{p}\overline{u}A(\overline{Δθ}_{N} − \overline{Δθ}_{S}) = 80 ± 37 MW (46%, ∼50%)

– p.13/19

Hydrography and heat flux above the ridge

\overline{H_{qs}} = 442 ± 213 MW. Max H_{qs} ≃ +2000 MW

– p.14/19

Lower heat flux budget:

H_{d} + H_{f} = H_{v} + H_{h}

H_{h} = 80 ± 37 MW, H_{v} = 640 ± 115 MW, H_{f} = 615 ± 120 MW H_{d} ≃ 105 MW versus 9000 ± 760 MW (Schultz et al, 1992) and ∼150 MW (Johnson et al, 2002) Partitioning of heat flux between sources: H_{d}:H_{f} ≃100:615≃1:6, rather than ∼10:1

Upper heat flux budget:

H_{v} = H_{p} ≃ \overline{H_{qs}}

Study  Estimate [MW] Heat flux
Stahr et al, 2003  640±115  H_{v}
Veirs et al, 2003  442±213  \overline{H_{qs}}
Thomson et al, 1992  2500±1525  H_{p}
Baker & Massoth, 1987  4250±2750  H_{p}
Rosenberg et al, 1988  3000±2000  H_{p}

– p.16/19

Conclusions

Net heat flux at MEF is ∼720 MW ∼10× geologic mean of ∼82 MW

MEF H_{d}:H_{f} ∼1:6, rather than 10:1

Expect multidirectional flow above ridge (high heat flux variance & Δx_{c} = 2.2 km)

When oscillatory currents disperse plumes, use a control volume to calculate net flux

Expect rectilinear tidal flow and mean inflows in valley (Sea Breeze hypothesis)

Acknowledgements

Russ McDuff Bill Lavelle Supervisory committee members Jeff Parsons Glenn Cannon Susan Hautala Will Wilcock Stephen Porter Family Annie Reese and Mila Val, Leslie, Laura, Pete Friends Many kind friends Fritz and Christian Graduate compatriots

– p.18/19

Russ McDuff Bill Lavelle Supervisory committee members Jeff Parsons Glenn Cannon Susan Hautala Will Wilcock Stephen Porter Family Annie Reese and Mila Val, Leslie, Laura, Pete Friends Many kind friends Fritz and Christian Graduate compatriots

– p.18/19

Russ McDuff Bill Lavelle Supervisory committee members Jeff Parsons Glenn Cannon Susan Hautala Will Wilcock Stephen Porter Family Annie Reese and Mila Val, Leslie, Laura, Pete Friends Many kind friends Fritz and Christian Graduate compatriots

– p.18/19

Russ McDuff Bill Lavelle Supervisory committee members Jeff Parsons Glenn Cannon Susan Hautala Will Wilcock Stephen Porter Family Annie Reese and Mila Val, Leslie, Laura, Pete Friends Many kind friends Fritz and Christian Graduate compatriots

lauraveirs.com

Russ McDuff Bill Lavelle Supervisory committee members Jeff Parsons Glenn Cannon Susan Hautala Will Wilcock Stephen Porter Family Annie Reese and Mila Val, Leslie, Laura, Pete Friends Many kind friends Fritz and Christian Graduate compatriots

– p.18/19

Questions?

Daniel McDuff, Whistler cornice (photo credit: Mark McDuff )

– p.19/19